OpenStudy (oaktree):

The mass of a cube of iron is 421 g. Iron has a density of 7.87 g/cm3. What is the mass of a cube of lead that has the same dimensions?

OpenStudy (anonymous):

multiply the mass of the cube by the reciprocal of the density (to get the units to cancel) and youre left with volume. multiply this volume by the density of lead (make sure the density has units that include cm^3) to get the mass of the cube of lead

OpenStudy (oaktree):

I've done that, but I'm apparently still wrong. Here's my work: 7.87 g/cm3=421cm3/V, therefore V=421/7.87 cm3 Density of lead=11.34 g/cm3 M=D*V=11.34*(421/7.87)=606.625... With significant figures, this is 607 g

OpenStudy (oaktree):

Can you tell me what I did? Because I can't find a mistake...

OpenStudy (anonymous):

i cant either, and i figured it out the same way you just did before you told me how you did it

OpenStudy (oaktree):

Well, I'm going nuts, because WebAssign is telling me it's wrong. Ah, well. I'll just ask my teacher. Thanks anyway.